# DeepLag: Discovering Deep Lagrangian Dynamics for Intuitive Fluid Prediction

### v2

@@ Abstract
Accurately predicting the future fluid is vital to extensive areas such as meteorology, oceanology, and aerodynamics. However, since the fluid is usually observed from an Eulerian perspective, its moving and intricate dynamics are seriously obscured and confounded in static grids, bringing thorny challenges to the prediction. This paper introduces a new Lagrangian-Eulerian combined paradigm to tackle the tanglesome fluid dynamics. Instead of solely predicting the future based on Eulerian observations, we propose DeepLag to discover hidden Lagrangian dynamics within the fluid by tracking the movements of adaptively sampled key particles. DeepLag utilizes the proposed where the Lagrangian movement of the tracked particles is inferred from Eulerian observations, and their accumulated Lagrangian dynamics information is incorporated into global Eulerian evolving features to guide future prediction respectively. Tracking key particles not only provides a transparent and interpretable clue for fluid dynamics but also makes our model free from modeling complex correlations among massive grids for better efficiency. Experimentally, DeepLag excels in three challenging fluid prediction tasks covering 2D and 3D, simulated and real-world fluids.
@@ 1 Introduction
Fluids, characterized by a molecular structure devoid of resistance to external shear forces, readily undergo deformation even under minimal force, thereby typically presenting extremely complex dynamics [10]. Consequently, the solvability of theorems for fluid mechanics, such as Navier-Stokes equations, is confined to a limited subset of flows due to the complex form and intricate multiphysics interactions [38]. In practical applications, computational fluid dynamics (CFD) is widely used to infer future fluid by numerical simulations. However, it suffers from substantial computation costs due to the curse of dimensionality [40]. Till now, accurately predicting future dynamics is still a challenging problem. Recently, benefiting from the tremendous non-linear modeling capability, deep models [8, 28, 22] have emerged as a promising tool for fluid prediction, which can be trained with CFD simulations or real-world observations and perform as a fast surrogate model during inference.
@@
A booming direction for deep fluid prediction is learning deep models to solve governing partial differential equations (PDEs) [45]. However, most of these methods [48, 29, 22, 21] attempt to capture fluid dynamics from the Eulerian perspective, which means modeling spatiotemporal correlations among massive grids remains unchanged over time. From this perspective, the complicated moving dynamics in fluids could be seriously obscured and confounded in static grids, bringing challenges in both computational efficiency and learning difficulties in accurately predicting future fluids.
@@
Apart from the Eulerian method, we notice another approach for elucidating fluid dynamics, the Lagrangian method [13], also known as the tracking method. This method primarily focuses on tracing individual fluid particles by conducting the temporal evolution of the position and velocity of each particle. Unlike the Eulerian methods, the Lagrangian approach describes the fluid dynamics through the moving trajectory of individual fluid particles, offering a more natural and neat representation of fluid dynamics with inherent advantages in capturing intricate flow dynamics. Moreover, it helps get around the Courant–Friedrichs–Lewy condition [6] for advection, allowing larger timesteps while retaining stability. As shown in Figure 1, we can find that fluid dynamics is much more apparent in the Lagrangian trajectory than the density change on static Eulerian grids.
@@
Building on the two perspectives mentioned earlier, we propose DeepLag as an Eulerian-Lagrangian Recurrent Network. Our aim is to integrate Lagrangian tracking into the deep model, thereby enhancing the dynamics modeling in Eulerian fluid prediction. To achieve this, we present the EuLag Block, a powerful tool that accomplishes Lagrangian tracking and Eulerian predicting at various scales. By leveraging the cross-attention mechanism, the EuLag Block assimilates tracked Lagrangian particle dynamics into the Eulerian field, guiding fluid prediction. It also forecasts the trajectory and dynamics of Lagrangian particles with the aid of Eulerian features. This unique Eulerian-Lagrangian design harnesses the dynamics information captured by Lagrangian trajectories and the fluid-structure feature learned in the Eulerian grid. In our experiments, DeepLag consistently outperforms existing models, demonstrating its state-of-the-art performance across three representative datasets, covering 2D and 3D fluid at various scales. Our contributions are as follows:
@@
Inspired by Lagrangian mechanics, we present EuLag Block, which can accurately track particle movements and interactively utilize the Eulerian feature and dynamic information in fluid prediction, enabling a better dynamic modeling paradigm.
@@ 2.1 Eulerian and Lagrangian methods
Eulerian and Lagrangian descriptions are two fundamental perspectives for describing fluid motion. The Eulerian view, commonly used in practical applications [23], observes fluid at fixed points and records physical quantities, such as density, as a function of position and time, . Thus, future fluid can be predicted by integrating velocity along the temporal dimension and interpolating results to observed grid points [49]. In contrast, the Lagrangian perspective focuses on the trajectory of individual particles, tracking a particular particle with initial position with its displacement at time . This approach reveals the intricate evolution of the fluid by following particle trajectories, making it convenient for describing complex phenomena like vortices, turbulence, and interface motions [39]. Two perspectives are constitutionally equivalent, as bridged by the velocity:
@@ 2.2 Neural Fluid Prediction
Neural PDE Solvers
This branch of methods in Figure 2(b) adopts deep models to learn the mapping from coordinates to correct values and formalize PDE constraints along with initial and boundary conditions as the loss function of deep models [48, 29, 46, 47]. Though this paradigm can precisely approximate the PDE solution, they usually require exact formalization for coefficients and conditions, limiting their generality and applicability to real-world fluids that are usually partially observed [35]. Plus, the Eulerian input disables them from handling Lagrangian descriptions.
@@
Recently, a new paradigm, illustrated in Figure 2(c), has emerged where deep models learn the neural operators between input and target functions, e.g., past observations to future fluid predictions. Various neural operators have significantly advanced fluid prediction following the DeepONet [22], which directly approximates mappings between equation parameters and solutions. For Eulerian grid data, models based on U-Net [32] and ResNet [15] structures have been proposed [30, 27, 17], as well as variants addressing issues like generalizing to unseen domains [44], irregular mesh [11], and uncertainty quantification [51]. Transformer-based models [43] enhance modeling capabilities and efficiency by exploiting techniques like Galerkin attention [3], incorporating ensemble information from the grid [14], applying low-rank decomposition to the attention mechanism [20], and leveraging spectral methods in the latent space [52]. Additionally, FNO [21] learns mappings in the frequency domain, and MP-PDE [2] utilizes the message-passing mechanism. For Lagrangian fluid particle data, some CNN-based methods [34, 42] model particle interactions through redesigned basic modules, while GNN-based methods [33, 25] update particle positions using the Encode-Process-Decode paradigm. Despite the progress made by these methods, they are limited to one description and do not organically combine both perspectives.
@@ 3 DeepLag
Following the convention [21], we formalize the fluid prediction problem as learning future fluid given past observation, as shown in Figure 2(d). Given a bounded open subset of -dimensional Euclidean space , a fluid with observed quantities can be seen as functions of coordinates in Banach space . and represent the fluid function at two consecutive time steps for the coordinate . Provided step observations , the fluid prediction process can be written as the following autoregressive paradigm:
where and represents the learned mapping. if and if .
@@ 3.1 Overall Framework
It is widely acknowledged that fluids exhibit different motion characteristics across varying scales [3, 52]. In order to capture the intrinsic dynamics information at different scales, we track the trajectories of the key particles on these scales separately and propose an Eulerian-Lagrangian Recurrent Network to realize the interaction between Eulerian and Lagrangian information.
@@
To better capture the complex dynamics within the fluid field, we sample crucial points from the given observations in the Eulerian perspective to determine the initial position for Lagrangian tracking in the dynamic sampling module. Explicitly, for the first predicting step, i.e. , given observations on all points with channels at the -th scale, we extract their spatial dynamics by a convolutional structure. We then calculate the probability matrix using softmax along spatial dimension:
where consists of a convolutional, activation, and fully-connected layer. We then sampled tracking particles using the probability matrix at the -th scale,
@@
For the -th step, we adopt a learnable embedding layer with multiscale architecture to encode past observations or predictions to obtain the Eulerian representations , where , and denotes the corresponding downsampled observation domain.
At each scale, we integrate EuLag Block to fuse the Lagrangian dynamics, thus directing the evolution of Eulerian features. This interaction also enables Eulerian features to guide the progression of Lagrangian dynamics. For the -th scale, we track key particles over time, with representing their positions and denoting their learned particle dynamics. As shown in Figure 2(d), the positions and dynamics of particles at the -th scale are learned autoregressively using the EuLag Block, which can be written as:
@@
After getting the output Eulerian features with Lagrangian particle position and dynamics , we further aggregated with at coarser scale by upsampling to be . Eventually, the prediction at step is decoded from with a projection layer. We unfold the implementation of the overall architecture in Appendix A.2.
@@ 3.2 EuLag Block
As formalized in Eq. (6), we adopt a recurrent framework to interactively exploit the information from Eulerian and Lagrangian perspectives, which consist of two main components: Lagrangian-guided feature evolving and Eulerian-conditioned particle tracking. For clarity, we omit the scale index in this section. The following quantities are all in the -th scale.
@@
where , and stand for linear projection layers. We summarize the above Lagrangian dynamics guided Eulerian Evolution process as .
@@ A.2.1 Sampling to initialize Lagrangian particles
Given the Eulerian input function at the -th scale, the operation in Eq. (4) is to extract the local dynamics around each Eulerian observation point with , and layers, which can be formalized as follows:
@@ A.2.2 Multiscale architecture
Given Eulerian fluid function at the -th time step, at previous time steps and the 0-1 boundary-geometry mask where 1 indicates the border and the unreachable area (like pillars in Bounded Navier-Stokes), the operation is to project original fluid properties in physical domain to deep representations with linear layer and position embedding, which can be formalized as follows:
@@ A.3 EuLag Block
where is described as Eq. (7). Moreover, we use the pre-normalization [53] technique for numerical stability.

### v3

@@ Abstract
Accurately predicting the future fluid is vital to extensive areas such as meteorology, oceanology, and aerodynamics. However, since the fluid is usually observed from the Eulerian perspective, its moving and intricate dynamics are seriously obscured and confounded in static grids, bringing thorny challenges to the prediction. This paper introduces a new Lagrangian-Eulerian combined paradigm to tackle the tanglesome fluid dynamics. Instead of solely predicting the future based on Eulerian observations, we propose DeepLag to discover hidden Lagrangian dynamics within the fluid by tracking the movements of adaptively sampled key particles. Further, DeepLag presents a new paradigm for fluid prediction, where the Lagrangian movement of the tracked particles is inferred from Eulerian observations, and their accumulated Lagrangian dynamics information is incorporated into global Eulerian evolving features to guide future prediction respectively. Tracking key particles not only provides a transparent and interpretable clue for fluid dynamics but also makes our model free from modeling complex correlations among massive grids for better efficiency. Experimentally, DeepLag excels in three challenging fluid prediction tasks covering 2D and 3D, simulated and real-world fluids.
@@ 1 Introduction
Fluids, characterized by a molecular structure devoid of resistance to external shear forces, readily undergo deformation even under minimal force, thereby typically presenting extremely complex dynamics [10]. Consequently, the solvability of theorems for fluid mechanics, such as Navier-Stokes equations, is confined to a limited subset of flows due to the complex form and intricate multiphysics interactions [38]. In practical applications, computational fluid dynamics (CFD) is widely used to infer future fluid by numerical simulations. However, it suffers from substantial computation costs due to the curse of dimensionality [40]. Accurately predicting future dynamics is still a challenging problem. Recently, benefiting from the tremendous non-linear modeling capability, deep models [8, 28, 22] have emerged as a promising tool for fluid prediction, which can be trained with CFD simulations or real-world observations and perform as a fast surrogate model during inference.
@@
A booming direction for deep fluid prediction is learning deep models to solve partial differential equations (PDEs) [45]. However, most of these methods [48, 29, 22, 21] attempt to capture fluid dynamics from the Eulerian perspective, which means modeling spatiotemporal correlations among massive grids unchanging over time. From this perspective, the complicated moving dynamics in fluids could be seriously obscured and confounded in static grids, bringing challenges in both computational efficiency and learning difficulties for accurately predicting future fluids.
@@
In parallel to the Eulerian method, we notice another major approach for elucidating fluid dynamics, the Lagrangian method [13], also known as the tracking method. This method primarily focuses on tracing individual fluid particles by conducting the temporal evolution of the position and velocity of each particle. Unlike the Eulerian methods, the Lagrangian approach describes the fluid dynamics through the moving trajectory of individual fluid particles, offering a more natural and neat representation of fluid dynamics with inherent advantages in capturing intricate flow dynamics. Moreover, it helps get around the Courant–Friedrichs–Lewy condition [6] for advection, allowing a larger inference step while retaining stability. In Figure 1, we can find that fluid dynamics is much more visually apparent in Lagrangian trajectories than density changes on static Eulerian grids.
@@
Building on the two perspectives mentioned earlier, we propose DeepLag as an Eulerian-Lagrangian Recurrent Network. Our aim is to integrate Lagrangian tracking into the deep model, thereby enhancing the dynamics modeling in Eulerian fluid prediction. To achieve this, we present the EuLag Block, a powerful module that accomplishes Lagrangian tracking and Eulerian predicting at various scales. By leveraging the cross-attention mechanism, the EuLag Block assimilates tracked Lagrangian particle dynamics into the Eulerian field, guiding fluid prediction. It also forecasts the trajectory and dynamics of Lagrangian particles with the aid of Eulerian features. This unique Eulerian-Lagrangian design harnesses the dynamics information captured by Lagrangian trajectories and the fluid-structure features learned in the Eulerian grid. In our experiments, DeepLag consistently outperforms existing models, demonstrating state-of-the-art performance across three representative datasets, covering 2D and 3D fluid at various scales. Our contributions are as follows:
@@
Inspired by Lagrangian mechanics, we present EuLag Block, which can accurately track particle movements and interactively utilize the Eulerian features and dynamic information in fluid prediction, enabling a better dynamic modeling paradigm.
@@ 2.1 Eulerian and Lagrangian Methods
Eulerian and Lagrangian descriptions are two fundamental perspectives for modeling fluid motion. The Eulerian view, commonly used in practical applications [23], observes fluid at fixed points and records physical quantities, such as density, as a function of position and time, . Thus, future fluid can be predicted by integrating velocity along the temporal dimension and interpolating the results to observed grid points [49]. In contrast, the Lagrangian view focuses on the trajectory of individual particles, tracking a particular particle from initial position by its displacement at time . This approach reveals the intricate evolution of the fluid by following particle trajectories, making it convenient for describing complex phenomena like vortices, turbulence, and interface motions [39]. Two perspectives are constitutionally equivalent, as bridged by the velocity:
@@ 2.2 Neural Fluid Prediction
Physics-informed Neural Networks (PINNs)
This branch of methods in Figure 2(b) adopts deep models to learn the mapping from coordinates to solutions and formalize PDE constraints along with initial and boundary conditions as the loss function [48, 29, 46, 47]. Though this paradigm can precisely approximate the PDE solution, they usually require exact formalization for coefficients and conditions, limiting their generality and applicability to real-world fluids that are usually partially observed [35]. Plus, the Eulerian input disables them from handling Lagrangian descriptions.
@@
Recently, a new paradigm, illustrated in Figure 2(c), has emerged where deep models learn the neural operators between input and target functions, e.g., past observations to future fluid predictions. Various neural operators have significantly advanced fluid prediction following the DeepONet [22], which directly approximates mappings between equation parameters and solutions. For Eulerian grid data, models based on U-Net [32] and ResNet [15] structures have been proposed [30, 27, 17], as well as variants addressing issues like generalizing to unseen domains [44], irregular mesh [11], and uncertainty quantification [51]. Transformer-based models [43] enhance modeling capabilities and efficiency by exploiting techniques like Galerkin attention [3], incorporating ensemble information from the grid [14], applying low-rank decomposition to the attention mechanism [20], and leveraging spectral methods in the latent space [52]. Additionally, FNO [21] learns mappings in the frequency domain, and MP-PDE [2] utilizes the message-passing mechanism. For Lagrangian fluid particle data, some CNN-based methods [34, 42] model particle interactions through redesigned basic modules, while GNN-based methods [33, 25] update particle positions using the Encode-Process-Decode paradigm. Despite the progress made by these methods, they are limited to one description and do not seamlessly combine Eulerian and Lagrangian views.
@@ 3 DeepLag
Following the convention [21], we formalize the fluid prediction problem as learning future fluid given past observation, as shown in Figure 2(d). Given a bounded open subset of -dimensional Euclidean space and an Eulerian space with observed physical quantities, and represent the Eulerian fluid field observation at two consecutive time steps on a finite coordinate set . Provided step observations , the fluid prediction process can be written as the following autoregressive paradigm:
where and represents the learned mapping between and the predicted field .
@@ 3.1 Overall Framework
It is widely acknowledged that fluids exhibit different motion characteristics across varying scales [3, 52]. In order to capture the intrinsic dynamics information at different scales, we track the trajectories of the key particles on scales separately and propose an Eulerian-Lagrangian Recurrent Network to realize the interaction between Eulerian and Lagrangian information. For clarity, we omit the scale index for primary quantities in this subsection, where .
@@
To better capture the complex dynamics within the fluid field, we sample crucial points from the given observations in the Eulerian perspective to determine the initial position for Lagrangian tracking in the dynamic sampling module. Explicitly, for the first predicting step, i.e. , given observations on all points in the observation domain of each scale, we extract their spatial dynamics by a convolutional network. We then calculate the probability matrix using softmax along the spatial dimension:
where consists of a convolutional, activation, and fully-connected layer. We then sample Lagrangian tracking particles using the probability matrix at each scale,
@@
For the -th step, we adopt a learnable embedding layer with multiscale downsampling to encode past observations or predictions to obtain the Eulerian representations at each scale.
Next, we integrate EuLag Block to fuse the Lagrangian dynamics at each scale, thus directing the evolution of Eulerian features. This interaction also enables Eulerian features to guide the progression of Lagrangian dynamics. For the -th scale, we track key particles over time, with representing their positions and denoting their learned particle dynamics. As shown in Figure 2(d), the positions and dynamics of particles at the -th scale are learned autoregressively using the EuLag Block, which can be written as:
@@
After getting the output Eulerian features with Lagrangian particle position and dynamics at the -th scale, we further aggregate with the predicted Eulerian field at coarser scale by upsampling . Eventually, the prediction at step is decoded from with a projection layer. We unfold the implementation of the overall architecture in Appendix A.2.
@@ 3.2 EuLag Block
As formalized in Eq. (6), we adopt a recurrent framework to interactively exploit the information from two fluid descriptions, which consist of two main components: Lagrangian-guided feature evolving and Eulerian-conditioned particle tracking. The following quantities are all in the -th scale.
@@
where , and stand for linear projection layers. We wrap the above attention into a standard transformer block with residual as Lagrangian-dynamics-guided Eulerian Evolution process: .
@@ A.2.1 Sampling to initialize Lagrangian particles
Given the Eulerian input field at the -th scale, the operation in Eq. (4) is to extract the local dynamics around each Eulerian observation point with , and layers, which can be formalized as follows:
@@ A.2.2 Multiscale architecture
Given Eulerian fluid observation at the -th time step, at previous time steps and the 0-1 boundary-geometry mask where 1 indicates the border and the unreachable area (like pillars in Bounded Navier-Stokes), the operation is to project original fluid properties in physical domain to deep representations with linear layer and position embedding, which can be formalized as follows:
@@ A.3 EuLag Block
where is described as Eq. (7), . Moreover, we use the pre-normalization [53] technique for numerical stability.